OpenStudy (anonymous):

integral of sqrt(1+cosx)

OpenStudy (anonymous):

\[\frac{ 2*\sin(x) }{ \sqrt{\cos(x)+1} }\]

OpenStudy (loser66):

@Spacelimbus I think we can convert the integrand by c0s(x/2) = sqrt (( 1+cosx))/2 what do you think?

OpenStudy (loser66):